Really just caught on these two:
Write the equation of the line which passes through (-4, 2) and is parallel to y = -x + 6 in slope-intercept form.
Write the equation of the line which passes through (2, -3) and is perpendicular to y = 4x + 7 in standard form.

Answer by Cromlix(4381)

y = -x + 6
Gradient = -1
Lines parallel to one another
have the same gradients (m)
m1 = m2
Using y - b = m(x - a) with Coords (-4,2)
y - 2 = -1(x -(-4)
y - 2 = -1(x + 4)
y - 2 = -x - 4
y = -x -4 + 2
y = -x - 2
y = 4x + 7
Gradient = 4
Lines perpendicular to one
another have gradients that
multiply together to give -1
m1 x m2 = -1
So second gradient = -1/4
4 x -1/4 = -1
Using y - b =m(x - a)with Coords (2,-3)
y - (-3) = -1/4(x - 2)
y + 3 = -1/4x + 1/2
y = -1/4x + 1/2 - 3
y = -1/4x - 5/2
OR
4y = -x - 10 (Multiplying by 4)
Hope this helps.
:-)